# DiffuGR: Generative Document Retrieval with Diffusion Language Models

### v4

@@ Abstract.
Generative retrieval (GR) re-frames document retrieval as a sequence-based document identifier (DocID) generation task, memorizing documents with model parameters and enabling end-to-end retrieval without explicit indexing. Existing GR methods are based on auto-regressive generative models, i.e., the token generation is performed from left to right. However, such auto-regressive methods suffer from: (1) mismatch between DocID generation and natural language generation, e.g., an incorrect DocID token generated in early left steps would lead to totally erroneous retrieval; and (2) failure to balance the trade-off between retrieval efficiency and accuracy dynamically, which is crucial for practical applications.
To address these limitations, we propose generative document retrieval with diffusion language models, dubbed DiffuGR. It models DocID generation as a discrete diffusion process: during training, DocIDs are corrupted through a stochastic masking process, and a diffusion language model is learned to recover them under a retrieval-aware objective. For inference, DiffuGR attempts to generate DocID tokens in parallel and refines them through a controllable number of denoising steps. In contrast to conventional left-to-right auto-regressive decoding, DiffuGR provides a novel mechanism to first generate more confident DocID tokens and refine the generation through diffusion-based denoising. Moreover, DiffuGR also offers explicit runtime control over the quality–latency tradeoff. Extensive experiments on benchmark retrieval datasets show that DiffuGR is competitive with strong auto-regressive generative retrievers, while offering flexible speed and accuracy tradeoffs through variable denoising budgets. Overall, our results indicate that non-autoregressive diffusion models are a practical and effective alternative for generative document retrieval. [^1] This paper is under review.
@@
Relevance Statement. This paper proposes a new diffusion-based generative document retrieval method DiffuGR. It aligns directly with the Web track topic of Search and retrieval-augmented AI and can be deployed in real-world Web systems, such as search engine, to obtain better search results.
@@ 1. Introduction
Generative Retrieval. Information retrieval (IR) is a fundamental component in modern artificial intelligence systems, powering search engines, recommendation systems, and large-scale knowledge access. In recent years, generative retrieval (GR) has emerged as a promising paradigm that replaces traditional index-based retrieval pipelines with an end-to-end generative model capable of directly producing document identifiers (DocIDs) in response to a query. GR first assigns each document with a unique sequence of tokens, aka DocID, then a generative sequence-to-sequence model is utilized to mapping the input sequence of a query to the output sequence of DocID tokens. By integrating retrieval tightly with generation and memorizing the document with model parameters, GR offers new potential to unify retrieval and reasoning, reduce system complexity, and unlock more flexible knowledge access.
@@
Figure 1. In contrast to auto-regressive generative retrieval (above), which generates the DocID from left to right, DiffuGR (below) efficiently generates multiple tokens in parallel and then re-masks uncertain tokens. The generation is progressively refined over multiple denoising steps.
@@
Auto-regressive Generation for GR. Most existing GR methods are built upon auto-regressive language models, which sequentially generate discrete identifier tokens in a left-to-right manner, as shown in the above part of Figure 1. Tay2022TransformerMA first proposed a differentiable search index (DSI) for GR, which utilizes a T5 language model as the generator. Plenty of research has been conducted following this research line, including constructing better DocIDs (Mehta2022DSIUT; sun2023learning; yang2023auto_search_indexer), using larger model parameters (pradeep2023doesscale; cai2025scalewang), and exploring new training and decoding strategies (sun2023learning; wu2025constrained).
Existing Limitations. Although these approaches have demonstrated effectiveness in achieving competitive performance compared to dense retrieval baselines, following limitations still exist:
(1) Mismatch between DocID generation and natural language generation. Different with natural language generation where the sequential order of words contains specific semantic information, tokens of a DocID may not have so strong left-to-right dependencies. It indicates that the auto-regressive left-to-right generation of DocID tokens could limit the retrieval performance. For example, if an incorrect DocID token is generated in early left steps, the following decoding generation steps would become meaningless since the correct DocID would never be generated.
(2) Failure to dynamically control the trade-off between efficiency and accuracy. Existing auto-regressive based GR methods can only generate DocID tokens one by one during inference with squared increasing computational cost. However, the capability to dynamically control the trade-off between inference efficiency and retrieval accuracy is of vital importance for real-world scenarios.
@@
The Proposed Method. To address these limitations, we proposed generative document retrieval with diffusion language models, dubbed DiffuGR, which models DocID generation as a discrete diffusion process. Specifically, for each document, DiffuGR first assigns a specific sequence of tokens as the DocID for this document. DiffuGR involves two kinds of DocIDs, including learnable DocIDs and linguistic DocIDs. Then, for model training, DocIDs are corrupted through a stochastic masking process, and a diffusion language model is learned to recover the masked tokens under a retrieval-aware objective function. At the inference stage, given an input query, DiffuGR generates DocID tokens in parallel and refines them through a controllable number of denoising steps, as shown in the bottom part of Figure 1.
@@
DiffuGR offers several appealing properties for GR compared to auto-regressive models. Firstly, its diffusion-based sampling enables parallel generation, where candidate DocID tokens are iteratively denoised rather than generated from left to right, allowing the model to revise and correct earlier errors instead of being constrained by erroneous token commitments. As a result, DiffuGR would first confirm more confident DocID tokens and gradually refine the generation across multiple denoising steps. Secondly, DiffuGR enables dynamic control of the trade-off between efficiency and accuracy through adjusting the number of denoising steps. For example, less denoising steps indicate faster inference speed while more denoising steps lead to higher retrieval accuracy. Besides, the stochastic nature of denoising allows DiffuGR to produce multiple plausible candidate DocIDs, thereby enhancing both the diversity and robustness of the retrieval results. Through these advantages, DiffuGR not only mitigates the generation mismatch but also opens up new possibilities for flexible quality–latency control.
Main Findings. In this paper, we conduct systematic study of DiffuGR. Our key findings are:
(1) Comparable performance. DiffuGR has achieved competitive retrieval performance on two widely used datasets, demonstrating the effectiveness of the proposed method.
@@
(2) Flexible runtime control. DiffuGR offers flexible runtime control over the quality-latency trade-off, which enables retrieval systems to perform dynamically scheduling based on load conditions.
(3) Existing weakness. Although experimental results have demonstrated the effectiveness of DiffuGR, the parallel generation indicates that DiffuGR does not support beam search, leading to trivial improvement on Recall@10 and Recall@100. It’s necessary to develop new methods for top-$k$ generation under diffusion scenarios.
Main Contributions. Our contributions can be summarized as:
  • We introduce diffusion language models as a new paradigm for GR, encouraging non-autoregressive DocID generation. To the best of our knowledge, this is the first work to utilize diffusion language models for end-to-end document retrieval.
  • We conduct extensive empirical comparisons between diffusion-based and auto-regressive GR, highlighting their respective strengths and weaknesses.
  • We demonstrate that DiffuGR not only matches the performance of auto-regressive GR methods, but also offers unique advantages of flexible quality-latency control.
To summarize, our work positions diffusion language models as a feasible and potentially superior alternative to auto-regressive models for GR. We hope this study opens up a new line of research at the intersection of generative modeling and information retrieval.
@@ 2. Related Work
In this section, we give a literature review regarding generative retrieval and diffusion models for text generation.
@@ 2.1. Generative Retrieval
Generative retrieval (GR) reformulates document retrieval as a generation task, in which a generative model (usually a seq2seq language model) directly generates the identifier of the target document. It provides an end-to-end solution for document retrieval and allows better utilization of large-scale generative language models. Document identifiers (DocIDs) are important for GR, through which retrieval systems can uniquely index the corresponding documents and enable end-to-end training and inference. Existing GR methods usually use learnable DocIDs or linguistic DocIDs.
Figure 2. An overview of DiffuGR, consisting DocID construction, model training, and inference. DiffuGR involves two kinds of DocIDs, i.e., linguistic DocIDs and learnable DocIDs. During training, the model is optimized to recover randomly masked DocID tokens. For inference, DiffuGR generates DocID tokens in parallel and refines them across multiple denoising steps.
Learnable DocIDs are usually constructed upon dense document representations, either through hierarchical $k$-means clustering (Tay2022TransformerMA; Wang2022ANC), product quantization (Zhou2022UltronAU; mekonnen2025lightweight), residual vector quantization (TIGER-Rajput; IRGen-Zhang; zeng2023scalable_and_effective), or progressively learning  (sun2023learning; yang2023auto_search_indexer). This process targets on compressing the semantic information contained in dense document representations into discrete tokens. Although promising results have been achieved, learnable DocIDs require expansion of the vocabulary, e.g., introducing additional trainable codebooks.
Linguistic DocIDs use strings, such as title, URL, or keywords, to represent documents. They naturally carry semantics related to associated documents, and the construction cost is relatively low, without the need for additional manual supervision. Plenty of work has been conducted based on linguistic DocIDs, yielding excellent results(DeCao2020AutoregressiveER; de-cao-etal-2022-multilingual; Bevilacqua2022AutoregressiveSE; Zhou2022UltronAU; Chen2022CorpusBrainPA; zhao2025unifying; UGR-Chen; tang2023inspired_by_learning_strategy; uni_gen; zhang2025replication).
Despite the success of GR, most existing methods generate DocID tokens in an auto-regressive left-to-right manner. Although qiao2023diffusionret proposed DiffusionRet which employs the diffusion model to generate a pseudo document for a query, the generation of DocIDs is still auto-regressive. However, such auto-regressive left-to-right generation of DocID tokens could limit the retrieval performance. Specifically, if an incorrect DocID token is generated in early left steps, the correct DocID would never be generated. Based on the above considerations, this work explores end-to-end generation of DocIDs through non-autoregressive diffusion language models (nie2025large; ye2025dream). Its core advantage lies in the fact that the diffusion model does not restrict the left-to-right generation order; instead, the model is encouraged to automatically select the most appropriate generation order based on its own conditions and confidence.
@@ 2.2. Diffusion Models for Text Generation
Diffusion models are initially introduced in continuous domains such as image generation (song2020denoising; ho2020denoising), and their success has inspired extensions to natural language processing by modeling text in continuous embedding spaces (li2022diffusion; gong2022diffuseq; gong-etal-2023-diffuseq). These continuous diffusion language models demonstrate the feasibility of applying iterative denoising for text generation, for both pre-training and fine-tuning on transformer architectures (genie2023). To better align with the discrete nature of text, discrete diffusion language models were later developed (austin2021structured; hoogeboom2021argmax; campbell2022continuous), which progressively corrupt token sequences and then reconstruct the original text during the reverse process. Further benefits have been observed through initializing from pre-trained masked language models such as BERT (he-etal-2023-diffusionbert; ye2025diffusionlanguagemodelsperform). Despite generation quality, recent work highlights the strengths of diffusion models in hybrid block-wise generation that balance sequential coherence with parallelism (ye2024diffusion; ye2024beyond; ye2025implicit; arriola2025blockdiffusion).
Scaling diffusion models to large language models (LLMs) has further extended their applicability. gulrajani2023likelihoodbased analyzed scaling laws for continuous diffusion.  lou2023discrete showed that discrete masked diffusion models have achieved perplexities competitive with GPT-2. Building on this, large-scale efforts such as DiffuGPT and DiffuLLaMA adopt pretrained auto-regressive LLMs into diffusion-based frameworks (gong2025scalingdiffusionlanguagemodels), while LLaDA (nie2025large) demonstrated that diffusion models trained from scratch at 8B parameters can rival strong auto-regressive LLMs like LLaMA3-8B. Recent commercial systems such as Mercury Coder (inceptionlabs_mercury) further validate the practicality of diffusion models for text and code generation.
In summary, these advances establish diffusion models as a viable alternative for natural language generation. While in this work, we explore the new potential to use diffusion models for end-to-end document retrieval, i.e., the generation of DocIDs for a query.
@@ 3. Method
In this section, we present details of the proposed DiffuGR. We first formulate the task of generative document retrieval in Section 3.1. The construction of DocIDs is illustrated in Section 3.2, then the training procedure of DiffuGR is described in Section 3.3. Finally, model inference is detailed in Section 3.4. Figure 2 shows an overall view of the proposed DiffuGR.
@@ 3.1. Task Formulation
This paper focuses on the task of generative retrieval, where the goal is to directly generate a document identifier (Docid) given a user query. Formally, let $\mathcal{D}=\{d_{1},d_{2},\dots,d_{N}\}$ denote the document collection, and $\mathcal{Q}=\{q_{1},q_{2},\dots\}$ denote user queries. Each document $d$ is associated with a sequence of discrete tokens (DocID) $\mathbf{z}=\{z_{1},z_{2},\dots,z_{l}\}$, where $l$ is the length of DocIDs. Given a query $q$, the model is required to generate the corresponding $\mathbf{z}$ that identifies the relevant document.
@@ 3.2. Construction of DocIDs
A key factor in GR is the design of DocIDs. Unlike dense retrieval methods, where continuous vectors implicitly encode document semantics, GR requires DocIDs to be discrete, compact, and decodable. To systematically study the effect of DocIDs for DiffuGR, we explore both learnable DocIDs and n-gram based linguistic DocIDs.
@@ 3.2.1. Learnable DocIDs
In this setting, DocIDs are learned through residual quantization over dense document embeddings. Let $\mathbf{d}$ denotes the embedding of document $d$. The initial residual vector is defined as $\mathbf{r}_{0}=\mathbf{d}$. The quantization is conducted recursively. At $i$-th step, we find the nearest neighbor within the $i$-th codebook $\mathbf{E}_{i}=\{\mathbf{e}_{i}^{k}\in\mathbb{R}^{d}\mid k=1,\dots,K_{i}\}$, where $K_{i}$ is the size of the $i$-th codebook, selecting the closest code embedding $\mathbf{e}_{i}^{z_{i}}$ to the current residual vector:
@@
The residual vector is updated as:
(2) $\mathbf{r}_{i}=\mathbf{r}_{i-1}-\mathbf{e}_{i}^{z_{i}},$
The original embedding $\mathbf{d}$ is approximated by adding the code embeddings. For training of the $i$-th codebook, we adopt the following loss:
@@
3.2.2. Linguistic DocIDs
In addition to learnable identifiers, we investigate n-gram based DocIDs, which are directly constructed from the text of each document. These DocIDs are lightweight, interpretable, and vocabulary-aligned, making them well-suited for language model generation. We consider two representative variants, including:
(1) Title-based DocIDs. The title of each document is directly adopted as the DocID for this document. Titles are typically concise human-written summaries of the main content, and hence naturally capture the document’s intent. However, not all documents have well-formed or unique titles. To address such issue, DiffuGR incorporates the second kind of DocIDs.

### v5

@@ Abstract.
Generative retrieval (GR) reframes document retrieval as an end-to-end task of generating sequential document identifiers (DocIDs). Existing GR methods predominantly rely on left-to-right auto-regressive decoding, which suffers from two fundamental limitations: (i) a mismatch between DocID generation and natural language generation, whereby an incorrect DocID token generated at an early step can lead to entirely erroneous retrieval; and (ii) an inability to dynamically balance the trade-off between retrieval efficiency and accuracy, which is crucial for practical applications.
To tackle these challenges, we propose generative document retrieval with diffusion language models, termed DiffuGR. DiffuGR formulates DocID generation as a discrete diffusion process. During training, DocIDs are corrupted through a stochastic masking process, and a diffusion language model is trained to recover them under a retrieval-aware objective. For inference, DiffuGR generates DocID tokens in parallel and refines them through a controllable number of denoising steps. Unlike auto-regressive decoding, DiffuGR introduce a novel mechanism to first generate plenty of confident DocID tokens and then refine the generation through diffusion-based denoising. Moreover, DiffuGR also offers explicit runtime control over the quality–latency tradeoff. Extensive experiments on widely-applied retrieval benchmarks show that DiffuGR outperforms strong auto-regressive generative retrievers. Additionally, we verify that DiffuGR achieves flexible control over the quality–latency trade-off via variable denoising budgets.
@@ 1. Introduction
Generative Retrieval. In recent years, generative retrieval (GR) has emerged as a promising paradigm that replaces traditional index-based retrieval pipelines with an end-to-end generative model capable of directly producing document identifiers (DocIDs) in response to a query (Cao et al., 2021; Tay et al., 2022; Wang et al., 2022b; Zhou et al., 2022b). By integrating retrieval tightly with generation and memorizing documents with model parameters, GR offers new potential to unify retrieval and reasoning, reduce system complexity, and unlock more flexible knowledge access.
@@
Figure 1. In contrast to auto-regressive generative retrieval (above), which generates the DocID from left to right, DiffuGR (below) efficiently generates multiple tokens in parallel and then re-masks uncertain tokens. The generation is progressively refined over controllable denoising steps.
@@
Most existing GR methods are built upon auto-regressive language models, which sequentially generate discrete identifier tokens in a left-to-right manner, as shown in the above part of Figure 1. Plenty of studies have been conducted following this research line, including constructing better DocIDs (Mehta et al., 2022; Sun et al., 2023; Yang et al., 2023; Zeng et al., 2024), using larger model parameters (Pradeep et al., 2023; Cai et al., 2025), and exploring new training and decoding strategies (Sun et al., 2023; Wu et al., 2025).
Limitations. Although these approaches have achieved competitive performance compared to dense retrieval baselines, several fundamental limitations still exist: (i) Mismatch between DocID generation and natural language generation. Unlike natural language generation where left-to-right word order encodes rich semantic, tokens of a DocID may not have so strong left-to-right dependencies. It indicates that the auto-regressive left-to-right generation of DocID tokens could limit the retrieval performance. For example, if an incorrect DocID token is generated in early left steps, the following decoding generation steps would become meaningless since the correct DocID would never be generated. (ii) Lack of dynamic control over the efficiency–accuracy trade-off. Existing auto-regressive based GR methods can only generate DocID tokens one by one during inference, leading to increasingly high computational costs as decoding progresses. More importantly, such approaches offer limited flexibility to adapt inference behavior to different operational requirements. In contrast, the ability to dynamically balance retrieval accuracy against inference efficiency is crucial for real-world applications with varied latency and resource constraints.
@@
The Proposed Method. To address the above challenges, we propose DiffuGR, a generative document retrieval framework based on diffusion language models, which formulates DocID generation as a discrete diffusion process. Specifically, DiffuGR assigns each document a sequence of tokens as its DocID. Two types of DocIDs are considered: learnable DocIDs and text-based DocIDs. During training, DocIDs are corrupted through a stochastic masking process, and a diffusion language model is learned to recover the masked tokens under a retrieval-aware objective function. At the inference stage, given an input query, DiffuGR generates DocID tokens in parallel and iteratively refines them through a controllable denoising steps, as shown in the bottom part of Figure 1.
@@
DiffuGR offers several appealing properties to GR: First, diffusion-based sampling enables parallel generation with iterative refinement: candidate DocID tokens are denoised jointly rather than generated sequentially. This allows the model to revise earlier predictions instead of being irrevocably constrained by early token errors. As a result, DiffuGR tends to establish high-confidence DocID tokens early and progressively refine the complete identifier across successive denoising steps. Second, DiffuGR provides dynamic control over the efficiency–accuracy trade-off by adjusting the number of denoising steps: fewer steps lead to faster inference, while additional steps improve retrieval accuracy. Finally, the stochastic nature of the denoising process enables DiffuGR to generate multiple plausible DocID candidates, thereby enhancing both the diversity and robustness of retrieval results.
@@
We conduct extensive experiments on NQ320K and MS MARCO datasets. Our key findings are twofold: (i) Improved effectiveness performance. DiffuGR consistently outperforms strong autoregressive generative retrievers on two widely used benchmark datasets, demonstrating the effectiveness of the proposed approach. (ii) Flexible runtime control. DiffuGR offers flexible runtime control over the quality-latency trade-off, which enables retrieval systems to perform dynamically scheduling based on load conditions.
Main Contributions. Our contributions can be summarized as follows: (1) We introduce diffusion language models as a new paradigm for GR, encouraging non-autoregressive DocID generation. To the best of our knowledge, this is the first work to utilize diffusion language models for end-to-end document retrieval. (2) We adapt the discrete diffusion paradigm to the generation of DocIDs. By leveraging the non-autoregressive nature and bidirectional context of diffusion models, our method mitigates the structural mismatch issue and alleviates the error accumulation problem inherent in auto-regressive GR. (3) We conduct extensive empirical comparisons between diffusion-based and auto-regressive GR, validating the effectiveness of DiffuGR. (4) We demonstrate that DiffuGR provides explicit and flexible control over the quality–latency trade-off, enabling dynamic adaptation of retrieval performance through adjustable denoising budgets.
Figure 2. An overview of DiffuGR, consisting DocID construction, model training, and inference. DiffuGR involves two kinds of DocIDs, i.e., text-based DocIDs and learnable DocIDs. During training, the model is optimized to recover randomly masked DocID tokens. For inference, DiffuGR generates DocID tokens in parallel and refines them across multiple denoising steps.
@@ 2.1. Generative Retrieval
Generative retrieval (GR) reformulates document retrieval as a generation task, in which a generative model (usually a seq2seq language model) directly generates the identifier of the target document. Existing GR methods usually use learnable DocIDs, text-based DocIDs, or number-based DocIDs. Learnable DocIDs are usually constructed upon dense document representations, either through hierarchical $k$-means clustering (Tay et al., 2022; Wang et al., 2022b), product quantization (Zhou et al., 2022b), residual vector quantization (Rajput et al., 2023; Zhang et al., 2023; Zeng et al., 2023), or progressively learning  (Sun et al., 2023; Yang et al., 2023). This process targets on compressing the semantic information contained in dense document representations into discrete tokens. Text-based DocIDs use strings, such as title, URL, or keywords, to represent documents. They naturally carry semantics related to associated documents, and the construction cost is relatively low, without the need for additional manual supervision. Number-based DocIDs use a single number or a series of numbers to represent documents (Tay et al., 2022; Zhou et al., 2022a). Although methods based on number-based DocIDs offer high retrieval efficiency, particularly when using single number, they are not considered in our experiments due to their poor generalization capability (Zhang et al., 2025). Plenty of work has been conducted based on text-based DocIDs, yielding excellent results(Cao et al., 2021; De Cao et al., 2022; Bevilacqua et al., 2022; Zhou et al., 2022b; Chen et al., 2022; Zhao et al., 2025; Chen et al., 2023; Tang et al., 2023; Li et al., 2024a; Zhang et al., 2025). Despite the success of GR, most existing methods generate DocID tokens in an auto-regressive left-to-right manner. Although Qiao et al. (2023) proposed DiffusionRet which employs the diffusion model to generate a pseudo document for a query, the generation of DocIDs are still auto-regressive. However, such auto-regressive left-to-right generation of DocID tokens could limit the retrieval performance. Specifically, if an incorrect DocID token is generated in early left steps, the correct DocID would never be generated. Based on above considerations, this work explores end-to-end generation of DocIDs through non-autoregressive diffusion language models (Nie et al., 2025; Ye et al., 2025).
@@ 2.2. Diffusion Models for Text Generation
Diffusion models are initially introduced in continuous domains such as image generation (Song et al., 2021; Ho et al., 2020), and their success has inspired extensions to natural language processing by modeling text in continuous embedding spaces (Li et al., 2022; Gong et al., 2023b, a). To better align with the discrete nature of text, discrete diffusion language models were later developed (Austin et al., 2021; Hoogeboom et al., 2021; Campbell et al., 2022), which progressively corrupt token sequences and then reconstruct the original text during the reverse process. Scaling diffusion models to large language models (LLMs) has further extended their applicability. Gulrajani and Hashimoto (2023) analyzed scaling laws for continuous diffusion.  Lou et al. (2024) showed that discrete masked diffusion models have achieved perplexities competitive with GPT-2. Building on this, large-scale efforts such as DiffuGPT and DiffuLLaMA adopt pretrained auto-regressive LLMs into diffusion-based frameworks (Gong et al., 2025), while LLaDA (Nie et al., 2025) demonstrated that diffusion models trained from scratch at 8B parameters can rival strong auto-regressive LLMs like LLaMA3-8B. While diffusion models have shown promise, effective constrained decoding remains challenging. Recently, Shi et al. (2025) proposed a specialized beam search for discrete diffusion in generative recommendation. In this work, we introduce and validate this decoding paradigm in the context of Generative Retrieval, demonstrating its capability to capture diverse retrieval candidates.
@@ 3. Method
We begin by formalizing generative document retrieval in Section 3.1. Section 3.2 then introduces the construction of DocIDs, followed by a description of the DiffuGR training in Section 3.3. Finally, we detail the inference process in Section 3.4. An overview of DiffuGR is illustrated in Figure 2.
@@ 3.1. Task Formulation
Generative retrieval aims to directly generate a document identifier (DocID) given a user query, where the generated DocID uniquely maps a corresponding document. Formally, let $\bm{D}=\{d_{1},d_{2},\dots,d_{N}\}$ denote a document collection, and $\bm{Q}=\{q_{1},q_{2},\dots,q_{M}\}$ denote user queries. Each document $d\in\bm{D}$ is associated with a unique identifier sequence (DocID), i.e., $\bm{z}=\{z_{1},z_{2},\dots,z_{l}\}$, where $l$ refers to the length of $\bm{z}$ and each token $z_{i}\in\bm{z}$ is drawn from residual quantization or token ID of the model vocabulary. Given a query $q_{i}\in\bm{Q}$, the model is required to generate the corresponding $\bm{z}$ that identifies the relevant document.
@@ 3.2. Construction of DocIDs
A key component of GR is the design of DocIDs. Unlike dense retrieval methods, GR requires DocIDs to be discrete, compact, and decodable. To systematically examine the role of DocID design in DiffuGR, we consider two classes of identifiers: learnable DocIDs and text-based DocIDs.
@@ 3.2.1. Learnable DocIDs
In this setting, DocIDs are learned through residual quantization over document embeddings. Let $\bm{d}$ denote the embedding of a document $d$ and $\bm{r}_{i}$ denote the residual vector of $i$-th step. The initial residual vector is set as $\bm{r}_{0}=\bm{d}$. Residual quantization is then performed recursively across multiple levels.
At $i$-th step, we assign $z_{i}$ as the index number of the closest code embedding in $i$-th codebook $\bm{E}_{i}=\{\bm{e}_{i}^{k}\in\mathbb{R}^{h}\mid k=1,\dots,K_{i}\}$ to $\bm{r}_{i-1}$, where $K_{i}$ is the size of $i$-th codebook and $\bm{r}_{i-1}$ denotes the residual vector from the previous step:
@@
then the $i$-th residual vector is updated as:
(2) $\bm{r}_{i}=\bm{r}_{i-1}-\bm{e}_{i}^{z_{i}}.$
The original embedding $\bm{d}$ is approximated by adding the code embeddings. For training of the $i$-th codebook, we adopt the following loss:
@@
3.2.2. Text-based DocIDs
In addition to learnable identifiers, we investigate text-based DocIDs, which are directly constructed from the text of each document. These DocIDs are lightweight, interpretable, and vocabulary-aligned, making them well-suited for language model generation. We consider two representative variants, including:
(1) Title-based DocIDs. The title of each document is directly adopted as the DocID for this document. Titles are typically concise human-written summaries of the main content, and hence naturally capture the document’s intent. However, not all documents have well-formed or unique titles. To address such issue, DiffuGR incorporates the second kind of text-based DocIDs.